Mickley Company’s plantwide predetermined overhead rate is $22.00 per direct labor-hour and its direct labor wage rate is $14.00 per hour. The following information pertains to Job A-500:

Direct materials $ 280
Direct labor $ 140

1. What is the total manufacturing cost assigned to Job A-500?

2. If Job A-500 consists of 30 units, what is the unit product cost for this job? (Round your answer to 2 decimal places.)
Answer #1

Solution 1:

Manufacturing cost assigned to Job A - 500 = Direct materials + Direct labor + Overhead applied

= $280 + $140 + ($140*22/14) = $640

Solution 2:

Unit product cost for this job = Manufacturing cost assigned to Job A / Nos of units

= $640 / 30 = $21.33 per unit